force Fi of magnitude 6.30 units acts at the origin in a direction 40.0 above the positive x axis. A second force F2 of magnitude 5.00 units acts at the origin in the direction of the positive y axis Find graphically the magnitude and direction of the resultant force + F2 units direction ecounterclockwise from the +x axis

A force  F1  of magnitude 6.30 units acts at the origin in a direction 40.0° above the positive x axis. A second force  F2  of magnitude 5.00 units acts at the origin in the direction of the positive y axis. Find graphically the magnitude and direction of the resultant force  F1 + F2.
Answer #1

Horizontal component, Fx = 6.3 cos40 = 4.83 unit

Vertical component, Fy = 6.3 sin40 + 5 = 9.05 units

Resultant force, F = sqrt(Fx^2 + Fy^2) = 10.26 units

Direction, theta = arctan(Fy/Fx) = 61.91 degree